Circular No. 6422

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996ai IN NGC 5005 Claudio Bottari, Sava, Italy, reports his discovery on a CCD image (+

ISIS deep-sky filter) obtained on June 16.86 UT of a supernova of mag 14.5 located some 50′′ east-northeast of the center of NGC 5005. S. Nakano, Sumoto, Japan, relays a subsequent unfiltered CCD observation by R. and Y. Kushida (Yatsugatake South Base Observatory) on June 19.472, which confirms the presence of a new star of mag 13.0 (near an apparent H II clump) that was not present on a May 12 image (limiting mag 16.5), located at α = 13^{h}10^{m}58^{s}.13, δ = +37^{o}03′35′′.4 (equinox 2000.0).

P. Garnavich, P. Challis, and R. Kirshner, Center for Astrophysics, communicate: “Spectra obtained by P. Challis with the Multiple Mirror Telescope and by P. Berlind with the 1.5-m Tillinghast telescope on June 20.3 UT show a red continuum with a strong, broad absorption feature at 614.81 nm. The spectra suggest that SN 1996ai is an extremely obscured type-Ia supernova, a few days before maximum light. A deep Na I interstellar absorption line is observed at 591.4 nm with an equivalent width of 0.23 nm. The steep red continuum implies E(B −V) > 1.0 when compared with SN 1996X at a similar phase. CCD images taken with the Whipple Observatory 1.2-m telescope by J. Kleyna on June 20.3 show the supernova 24′′ east and 4′′ north of the galaxy nucleus at V = 16.2 ± 0.2.”

QX NORMAE S. Wachter, E. W. Deutsch, J. D. Hughes, and B. Margon, University of

Washington; and A. Layden, McMaster University, report: “Observations made with the Cerro Tololo 0.9-m telescope of the transient x-ray source 4U 1608–52 since the recent x-ray outburst (IAUC 6331, 6336) show that the faint optical counterpart QX Nor (Grindlay and Liller 1978, Ap.J. 220, L127) has reappeared. Data obtained in May and June show the object at I = 20, ∼ 2 mag fainter than during the 1977 appearance. However, these recent observations are 2–3 months after the x-ray outburst, whereas the 1977 counterpart was observed only 1 month later; thus a similar maximum optical brightness at both x-ray outbursts is plausible. Even the faint current state is probably still significantly brighter than the true quiescent magnitude of the counterpart, which has not yet been established: a comparison of R images of the field obtained this year with identical data last year, when there is no optical detection (and presumably no x-ray activity), shows that the object has brightened by at least 2 mag in that 1-yr interval. A search for short-term variability in ∼ 10 nights of data is underway.”

1996 June 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green